OpenStudy (anonymous):

How do i write this as a vertex form f(x)=-7x^2+21x-9? when the vertex is -3/2 & -99/4? please help

OpenStudy (anonymous):

the vertex of the function you have is not \[(-\frac{3}{2},-\frac{99}{4})\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

it is \[(\frac{3}{2},\frac{27}{4})\]

OpenStudy (anonymous):

thats h & k

OpenStudy (anonymous):

oh ok. so how do i put it as a vertex form